Queen Mary Company manufactures and sells a telephone answering machine. The company’s contribution format profit and loss account for the most recent year is given below:
                                                              Total        Per unit      % of sales
Sales (20,000 units)                          ₤1,200,000        ₤60          100%
Less variable expenses                           900,000       ₤45            ?%
Contribution margin                               300,000       ₤15            ?%
Less fixed expenses                                240,000
Profit                                                      ₤60,000
Management is anxious to improve the company’s profit performance and has asked for several items of information.
Required:
1. Refer to the original data. Compute the company’s margin of safety in both pound and percentage form.
2. In an effort to increase sales and profits, management is considering the use of a higher-quality speaker. The higher-quality speaker would increase variable costs by ₤3 per unit, but management could eliminate one quality inspector who is paid a salary of ₤30,000 per year. The sales manager estimates that the higher-quality speaker would increase annual sales by at least 20%
(a) Assuming that changes are made as described above, prepare a projected profit and loss account for next year. Show data on a total, per unit, and percentage basis.
(b) Compute the company’s new break-even point in both units and pounds of sales. Use the CM method.
(c) Would you recommend that the changes be made?
Answer #1

1)the sales above break even sales is known as margin of safety

So margin of safety = sales - breakeven sales

= 20000- fixed cost / contribution per unit (BEP formulae)

= 20000- 240000/15= 4000units

Margin of safety in pound is 4000×20= 80000

As percentage of sales = 4000/20000×100= 20%

after doing the above changes

new sales is 24000(20000+20%)

new variable cost 48 per unit

new contribution per unit is 210000

2(a)

Projected profit and loss account

sales = 24000×60= 1440000(100%)

less variable expense = 48×24000= 1152000(80%)

contribution margin Is 12×24000=288000(20%)

less fixed cost 210000

profit is . 78000(5.4%of sales)

2(b)

new bep is = new fixed cost/ new contribution per unit

= 210000/12= 17500units

Bep in pounds is 17500× 60= 1050000

C)even though break even point increased after the given effects increase in sales has covered the loss and profit has been increased by18000 pounds so it is suggested for the company to make the changes